# AdaGC: Improving Training Stability for Large Language Model Pretraining

### v1

@@ 4.1 Preliminaries
$\theta_{t}=\theta_{t-1}-\alpha\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
To prevent unstable updates from gradient explosions, gradient clipping (Pascanu et al., 2013) modifies the update rule as:
$\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\alpha h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.99994pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
Here $\lambda_{\text{abs}}$ is an absolute clipping threshold requiring careful tuning. Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{\text{abs}}$) rather than value-based clipping (element-wise truncation).
4.2 Adaptive Gradient Clipping based on Local Gradient Norm
This section introduces a novel gradient clipping strategy termed Adaptive Gradient Clipping (AdaGC), which distinguishes itself by not relying on a global gradient norm. Instead, AdaGC focuses on the local gradient norm of each parameter and utilizes a dynamic adaptive mechanism for gradient clipping. The proposed method employs an Exponential Moving Average (EMA) mechanism to maintain smoothed estimates of historical gradient norms per parameter, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each parameter’s specific conditions. EMA is widely used in deep learning, and within AdaGC, it facilitates the balancing of historical and current gradient norms. The formulation is as follows:
$\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[10.00002pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th parameter at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$. Additionally, $\beta$ is the smoothing coefficient for EMA. We consistently incorporate the clipped gradient norm into the historical observations rather than the pre-clipped values.
Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each parameter’s gradient norm. Whenever the gradient norm at a current timestep exceeds a predefined range of average norms within a historical window, it effectively suppresses these outlier gradients.
However, during the initial stages of model training (e.g., the first 100 steps), the gradient norms are typically large and fluctuate significantly, indicating a substantial decreasing trend. Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to traditional global-norm-based methods, AdaGC might delay clipping, thus potentially slowing down the loss reduction. To address these issues, we introduce a hyperparameter $T_{start}$(default set to 100), representing a warm-up period during which traditional global gradient norm clipping is applied.
Additionally, the AdaGC strategy can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), enhancing its practicality and flexibility. Algorithm 1 demonstrates its implementation with the AdamW optimizer.
Overall, AdaGC offers a more precise and effective gradient management method for large-scale model training, contributing to improved training stability and performance.
@@ 4.2 Adaptive Gradient Clipping based on Local Gradient Norm
1:  given: $\{\alpha_{t}\}_{t=1}^{T},\lambda_{w},\epsilon_{1},\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}\in[0,1),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
@@
6:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}}$
7:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}}$
@@
9:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
@@
15:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
@@
18:    $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}$
19:    $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}^{2}$
@@
21:    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\alpha_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\alpha_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon_{1})$
@@
4.3 Convergence Analysis
In this section, we give the convergence guarantee of Adam with AdaGC stated as follows:
Theorem 4.1.

### v2

@@ 5.1 Experimental Setup
Comparison Methods. We focus on clipping-based methods and compare gradient and update clipping baselines, including GlobalGC (Pascanu et al., 2013), Gradient Value Clipping (ClipByValue), AGC (Brock et al., 2021), and Clippy (Tang et al., 2023). We also evaluate recent methods, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024). Results are in Appendix D.2 Table 12.
Training Details. Pre-training large-scale models is typically resource-intensive. In this study, our primary focus was to investigate training instability rather than to pursue ultimate accuracy. To facilitate multiple experimental runs, we conducted 9,000 training steps over 36 billion tokens for both Llama-2 1.3B and 7B, 36,000 steps over 36 billion tokens for Mixtral 8x1B, and 60,000 steps over 1 trillion tokens for ERNIE 10B-A1.4B, the latter additionally serving to validate the long-term stability of AdaGC. Further details regarding hyperparameters can be found in Table 9 in Appendix B.
Evaluation Metrics. To quantitatively assess training stability, we follow (OLMo et al., 2024; Karpathy, 2024) and adopt the spike score as an objective metric. Specifically, the spike score is defined as the percentage of values in a time series that deviate by at least ten standard deviations from a rolling average of the preceding 1,000 values. This metric is primarily applied to training loss to detect sudden instabilities. Additionally, we evaluate performance using the training loss and validation perplexity (PPL) curves, as well as standard benchmark results, to provide a comprehensive assessment of convergence efficiency and model quality.
Standard Benchmark. We conducted a comprehensive evaluation of the model’s zero-shot and two-shot capabilities across seven well-established benchmarks: ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), and MMLU (Hendrycks et al., 2020). Following standard practice (Zhang et al., 2025), we report accuracy norm for ARC-E, ARC-C, HellaSwag, OBQA, and SciQ, as well as standard accuracy for all other tasks. For ERNIE 10B-A1.4B, which has been trained on 1T tokens, we evaluate its general abilities on a range of benchmarks, including MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2021), and HumanEval (Chen et al., 2021). These benchmarks assess the model’s enhanced capabilities in performing diverse downstream tasks, such as examination, reasoning, factuality, and coding.
(a) Llama-2 7B training dynamics.
(b) Mixtral 8x1B training dynamics.
Figure 3: Large language model training analysis: Llama-2 7B and Mixtral 8x1B model comparison shows AdaGC’s loss spike elimination and performance gains.
5.2 Critical Hyperparameter Selection
We systematically evaluated two key hyperparameters in AdaGC: the EMA coefficient $\beta$ and the relative clipping threshold $\lambda_{rel}$. Specifically, we performed a grid search on the Llama-2 7B model to optimize these two hyperparameters, using zero-shot and two-shot performance across multiple tasks as evaluation metrics. As shown in Tables 4 and 4, the best performance was achieved when $\lambda_{rel}=1.04$ and $\beta=0.99$. We therefore adopted this configuration as the default setting for subsequent experiments and terminated further hyperparameter search. In addition, as observed in Tables 4 and 4, AdaGC’s performance remains relatively stable across different hyperparameter values, suggesting that the method is robust to hyperparameter variations.
5.3 Main Experimental Results
[…]
Table 5: The Zero-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B GlobalGC 43.18 25.68 57.19 46.62 30.20 69.97 52.64 22.97 68.40 46.32
ClipByValue 42.17 25.68 59.94 44.11 30.40 69.59 53.28 22.99 68.00 46.24
Clippy 41.71 24.66 56.51 45.43 30.00 69.21 54.85 22.90 67.50 45.86
AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
Llama-2 7B GlobalGC 49.49 27.56 56.30 56.06 33.60 74.59 55.33 23.12 71.20 49.69
ClipByValue 46.21 26.88 57.03 53.49 33.20 71.65 53.59 23.36 70.50 48.43
AGC 48.15 28.16 52.87 55.47 32.80 72.74 57.85 24.33 71.70 49.34
Clippy 47.69 27.73 57.46 53.34 32.40 72.74 54.38 25.36 73.40 49.39
AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
Mixtral 8x1B GlobalGC 44.70 25.94 56.57 53.08 33.00 71.60 54.70 22.91 67.20 47.74
AdaGC 46.68 26.37 58.93 55.85 32.20 73.12 54.38 23.22 70.30 49.01
Table 6: Evaluation results of ERNIE 10B-A1.4B on multiple benchmarks after 21,000 (350B tokens) and 60,000 (1T tokens) training steps, comparing different optimization configurations.
Steps (tokens) Method AdamW eps MMLU GSM8K BBH TruthfulQA HumanEval Avg.
21k (350B) GlobalGC 1e-8 47.75 28.35 28.80 22.02 19.51 28.09
GlobalGC 1e-15 39.11 21.46 29.35 23.39 15.24 25.71
AdaGC 1e-15 42.07 25.32 27.89 24.92 20.73 28.19
60k (1T) GlobalGC 1e-8 48.61 39.88 30.84 30.73 22.56 34.52
GlobalGC 1e-15 48.48 40.79 30.59 28.29 23.78 34.38
AdaGC 1e-15 48.70 36.01 31.38 35.02 22.56 34.73
[…]
Table 7: Comparison of spike scores for various models under different clipping methods.
Model Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
Method GlobalGC ClipByValue AGC Clippy AdaGC GlobalGC AdaGC GlobalGC AdaGC
Total Steps 9K 9K 9K 9K 9K 36K 36K 21K 21K
Num Spikes 3 9 8 3 0 52 0 2 0
Spike Score (%) 0.0333 0.1000 0.0889 0.0333 0.0000 0.0144 0.0000 0.0100 0.0000
Table 8: GPU hours under the same configuration. DPS denotes distributed parallel strategies.
Model Llama-2 1.3B Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
DPS DP=256, TP=1, PP=1 DP=32, TP=2, PP=1 DP=256, TP=1, PP=1, EP=1 DP=64, TP=1, PP=4, EP=8
Steps 9K 9K 36K 21K
GlobalGC 513.0 1468.2 2060.8 22922
AdaGC 511.6 1402.4 2035.2 22572
[…]
5.4 Optimizer Compatibility: Muon and Lion
AdaGC is an optimizer-agnostic gradient clipping method that can be seamlessly integrated not only with AdamW, but also with other optimizers. To verify the generality of AdaGC, we conducted experiments on both LLM and VLM tasks by combining Llama-2 1.3B and CLIP ViT-Base models with the Muon and Lion optimizers, respectively, and compared them against GlobalGC. Although no loss spikes were observed under either experimental setting, AdaGC consistently demonstrated strong compatibility and generalization. In downstream zero-shot average accuracy, AdaGC outperformed GlobalGC by 0.14% (47.18% vs. 47.04%) with Muon and by 0.16% (40.81% vs. 40.65%) with Lion. These results further confirm that AdaGC can be effectively applied across different optimizers, providing stable training and improved downstream performance.
5.5 End-to-End training Wall-clock
Table 8 compares the GPU hours required for training various models using different distributed parallelism strategies. Compared to GlobalGC, AdaGC reduces end-to-end GPU hours by 0.27% on Llama-2 1.3B, 4.48% on Llama-2 7B, 1.24% on Mixtral 8x1B, and 1.53% on ERNIE 10B-A1.4B, thanks to reduced communication overhead. This highlights AdaGC’s additional communication and efficiency benefits in large-scale distributed training.
5.6 Ablation Study
We conduct systematic ablation studies across three critical dimensions of AdaGC: (1) EMA gradient norm initialization strategies, (2) GlobalGC warm-up steps, (3) adaptivity efficacy, and (4) locality granularity.
(a) Training dynamics.
(b) Average accuracy.
Figure 4: AdaGC with Muon on Llama-2 1.3B.
(c) Training dynamics.
(d) Average accuracy.
Figure 5: AdaGC with Lion on CLIP ViT-Base.
(a) $\gamma_{t,i}$ initialization.
(b) $T_{start}$ warm-up.
(c) Adaptivity, locality.
Figure 6: Training dynamics of AdaGC ablations, showing (a) the influence of different EMA initialization strategies; (b) the impact of the GlobalGC warm-up steps $T_{start}$; and (c) the effects of adaptivity and locality granularity on gradient clipping efficacy and final loss.
EMA Initialization Strategy. The initialization of EMA gradient norms requires careful design due to large initial gradient fluctuations during early training phases (first 100 steps). We evaluate five initialization variants: The default AdaGC strategy employs GlobalGC during warm-up while tracking minimum per-parameter norms ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$). Comparative approaches include: (1) norm initialization without GlobalGC warm-up (directly using $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$ from step 0), (2) constant initialization ($\gamma_{0,i}\in\{0.5,1.0\}$), and (3) thresholded initialization ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,0.1)$). Figure 6(a) demonstrates that while all variants eliminate loss spikes, convergence quality varies within 0.36%. The default strategy achieves optimal final loss (2.9708 vs 2.9725 for next-best), showing that GlobalGC-guided warm-up better preserves parameter update consistency than direct initialization. This establishes the importance of phased initialization for gradient norm adaptation.
Warm-up Steps $T_{start}$. To further investigate whether the choice of GlobalGC warm-up steps $T_{start}$ has a significant impact and to provide practical guidance for practitioners, we additionally evaluate $T_{start}=\{0,50,100,150,200,500,1000,2000\}$. The results in Figure 6(b) show that $T_{start}=100$ consistently achieves the best performance. According to the EMA initialization formula $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$, an excessively large $T_{start}$ accumulates lower $\gamma_{t,i}$ values due to early training dynamics, which may lead to over-clipping and suppressed convergence in later training. Conversely, an overly small $T_{start}$ accumulates larger $\gamma_{t,i}$ values, which may delay clipping and hinder timely suppression of abnormal gradients. In contrast, $T_{start}=100$ introduces negligible additional overhead for large-scale training while providing consistently stable performance improvements.
Adaptivity Efficacy and Locality Granularity. We conduct three sets of ablation experiments to evaluate the adaptivity and locality of AdaGC. The baseline uses GlobalGC (no adaptivity, no locality) with a fixed threshold of 1.0. In comparison, we examine (1) adaptive global gradient norm clipping (Global AdaGC, adaptive but non-local), which employs a single adaptive threshold for the entire model, (2) tensor-wise gradient norm clipping (TensorWiseGC, local but non-adaptive), which allocated each tensor’s fixed clipping threshold proportionally to its parameter count relative to the entire model, and (3) tensor-wise adaptation (AdaGC, adaptive and local), which adjusts thresholds independently for each tensor. As shown in Figure 6(c), Global AdaGC reduces but does not completely eliminate spike events (1 event vs. 0 for tensor-wise) and yields a 0.25% higher final loss (2.9639 vs. 2.9712). Although TensorWiseGC also mitigates loss spikes, it noticeably slows down convergence and requires careful per-tensor threshold tuning to perform well. These results demonstrate that tensor-wise adaptive clipping provides both greater spike suppression and lower loss than other approaches.
6 Conclusion
The factors triggering loss spikes in large-scale pretraining are diverse and remain an open research problem, with no unified solution to date. Unlike prior work that seeks to identify root causes, we focus on a gradient-centric remedy and introduce AdaGC, an adaptive per-tensor gradient clipping method that prevents abnormal gradients from contaminating optimizer states. This approach ensures smoother updates and effectively eliminates loss spikes. Extensive experiments demonstrate that AdaGC delivers robust and stable training across both dense and MoE models, from 1.3B to 10B parameters, consistently reducing spike scores to zero and improving benchmark performance. Our results highlight AdaGC as a simple and effective solution for stable large-scale model pretraining.
Impact Statement
Appendix Contents
(Run LaTeX twice to generate Appendix Contents.)
Appendix A Pseudocode of AdamW with AdaGC
Algorithm 1 presents the pseudocode of AdamW integrated with AdaGC. For clearer exposition, we highlight different components according to their origins: orange indicates the procedures inherited from the original GlobalGC algorithm, while blue is used to denote the new contributions and modifications introduced by AdaGC. Specifically, the GlobalGC steps include the global gradient clipping implemented via the scaling factor and the use of the clipped gradient in subsequent moments. The AdaGC components mainly comprise adaptive per-parameter clipping, the initialization and update of the adaptive threshold $\gamma_{t,i}$, and the warm-up strategy governed by $T_{start}$.
@@ Appendix A Pseudocode of AdamW with AdaGC
1:  given: $\{\eta_{t}\}_{t=1}^{T},\lambda_{w},\epsilon,\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta\in[0,1)},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
@@
6:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}$}
7:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}$}
@@
9:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
@@
15:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
@@
18:  $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}}$
19:  $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}^{2}}$
@@
21:  $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\eta_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon)$
@@
Appendix B Hyper-Parameters
B.1 Model Hyper-Parameters
Table 9 summarizes the model hyper-parameters used for all experiments. For each model, we report the core architecture settings (such as number of layers, hidden dimension, attention heads, and feedforward dimension), MoE-related configurations, and main optimization hyper-parameters (including learning rate, warmup, weight decay, and Adam parameters). Clipping thresholds $\lambda_{abs}$, $\lambda_{rel}$, and momentum $\beta$ are also listed, in correspondence with the techniques discussed in the main text. All experiments use a batch size and sequence length as shown, and we employ bfloat16 precision for most models except ERNIE, which uses float8. The symbol ‘–’ indicates settings not applicable to a specific architecture.
Table 9: Hyper-parameters used in our LLMs experiments. $\lambda_{abs}$ represents the absolute global clipping threshold of GlobalGC. $\lambda_{rel}$ and $\beta$ represent the relative clipping threshold and the momentum of our AdaGC, respectively. The symbol ‘–’ indicates that the parameter is not applicable.
Model Llama-2 1.3B Llama-2 7B ERNIE 10B-A1.4B Mixtral 8x1B
Precision bfloat16 bfloat16 float8 bfloat16
Num layers 24 32 25 24
Hidden dim size 2048 4096 2560 2048
FFN dim size 5461 11008 1024 5632
Num attention heads 32 32 20 32
Num key value heads 32 32 4 4
Attention bias
Num shared experts - - 1 0
Num router experts - - 48 8
Num experts per token - - 3 2
Sequence length 2048 2048 4096 2048
Batch size 2048 2048 4096 512
Iterations 9000 9000 21000 36000
LR decay cosine cosine wsd cosine
Warmup iterations 2000 2000 2000 500
Weight decay 0.1 0.1 0.1 0.1
Adam $\beta_{1}$ 0.90 0.90 0.90 0.90
Adam $\beta_{2}$ 0.95 0.95 0.95 0.999
$\lambda_{abs}$ 1.0 1.0 1.0 1.0
$\lambda_{rel}$ 1.04 1.04 1.04 1.04
$\beta$ 0.99 0.99 0.99 0.99
B.2 Clipping Hyper-Parameters
For other clipping methods, we primarily followed the recommended default settings from prior work, and performed limited tuning only when necessary to ensure a fair comparison.
Specifically:
  • GlobalGC: We used the commonly adopted global clipping threshold $\lambda_{abs}=1.0$ in large-scale pretraining.
  • ClipByValue: Following the SPAM (Huang et al., 2025) setting, we set the clipping threshold to $\lambda_{abs}=1e-3$.
  • AGC: We performed small-range tuning over $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to find the best setting.
  • Clippy: We tuned over $\lambda_{abs}\in\{0.1,0.3,0.5\}$ and $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to select the optimal combination.
  • SPAM: We adopted the default hyperparameters recommended for standard pretraining in the original paper, which were reported to perform well across diverse settings. Specifically, we set the interval to $\Delta T=500$, the threshold to $\theta=5000$, and the warmup steps to $N=150$.
The final hyper-parameters used for other clipping methods are summarized in Table 10.
Table 10: Hyper-parameters for other clipping methods.
Method Hyperparameters
GlobalGC $\lambda_{abs}=1.0$
ClipByValue $\lambda_{abs}=1e-3$
AGC $\lambda_{rel}=1e-3$
Clippy $\lambda_{abs}=0.5,\lambda_{rel}=1e-2$
SPAM $\Delta T=500$, $\theta=5000$, $N=150$
Appendix C Experimental Details for CLIP
To further investigate the optimizer compatibility of AdaGC, we evaluated its effect on large-scale vision-language model pre-training, focusing on the CLIP ViT-Base model (Radford et al., 2021) with the Lion optimizer (Chen et al., 2024). The model comprises 151 million parameters and is trained on the LAION-400M (Schuhmann et al., 2021) dataset. Training is conducted for 20,000 steps, covering 320M image-text pairs.
The key training hyper-parameters are as follows: a learning rate of 0.002, weight decay of 0.2, and batch size of 32,768. We employ patch-dropout with a drop rate of 0.5 (Li et al., 2023), following recent best practices (Wortsman et al., 2023). The learning rate is linearly warmed up for the first 5,000 steps (Goyal et al., 2017), and subsequently decayed according to a cosine schedule (Loshchilov & Hutter, 2016).
Following pre-training, we report downstream zero-shot evaluation results on the ImageNet (Russakovsky et al., 2015) validation set. The results are shown in Figure 5 in the main text.
Appendix D More Evaluation Results
D.1 Results on Downstream Benchmarks
The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks are presented in Table 11.
Table 11: The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks. The best scores in each column are bolded. HellaSw. = HellaSwag, W.G. = WinoGrande.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B GlobalGC 47.26 25.60 50.31 46.44 32.20 69.64 52.33 25.07 77.80 47.41
ClipByValue 47.10 25.77 56.54 43.97 30.00 68.88 52.96 26.09 77.20 47.61
Clippy 46.55 25.85 49.76 45.71 30.00 70.02 53.20 25.69 77.70 47.16
AdaGC 46.04 26.19 49.72 47.51 31.00 69.70 54.38 24.98 78.50 47.56
Llama-2 7B GlobalGC 55.81 28.58 60.70 56.54 33.00 73.72 56.75 25.51 83.20 52.64
ClipByValue 51.94 26.88 57.55 53.36 32.40 72.31 54.14 26.63 81.60 50.75
AGC 52.95 28.67 56.15 55.69 35.40 73.07 56.43 26.88 82.80 52.00
Clippy 52.86 29.10 56.48 53.76 31.80 73.07 55.72 26.03 82.60 51.27
AdaGC 56.86 29.61 59.36 57.89 33.60 73.99 57.62 26.46 85.90 53.47
Mixtral 8x1B GlobalGC 50.34 27.39 58.81 52.96 34.20 71.16 54.06 25.37 79.90 50.47
AdaGC 53.83 28.42 58.69 55.66 33.80 73.07 54.14 25.12 81.80 51.61
D.2 Results of Other Baseline Methods
Table 12: The Zero-Shot evaluation results of Llama-2 1.3B/7B models on standard benchmarks.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B WeSaR-GlobalGC 43.56 25.17 59.94 45.08 30.00 70.29 52.96 22.90 65.80 46.19
SPAM 42.05 24.83 59.60 42.82 30.00 69.31 52.17 23.02 66.40 45.58
ScaledEmbed-GlobalGC 42.21 25.51 59.66 45.50 31.80 70.02 53.28 23.22 65.20 46.27
AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
Llama-2 7B WeSaR-GlobalGC 49.75 27.22 56.12 55.38 33.80 73.39 56.27 23.02 71.40 49.59
SPAM 48.53 25.77 60.34 51.89 32.60 72.03 54.54 22.95 71.00 48.85
ScaledEmbed-GlobalGC 48.57 26.71 60.89 54.32 32.60 72.25 55.33 23.66 70.50 49.42
AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
Table 13: Comparison of spike scores for various models under different methods.
Model Llama-2 1.3B Llama-2 7B
Method WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC
Total Steps 9K 9K 9K 9k 9K 9K 9K 9K
Num Spikes 2 0 10 0 1 3 8 0
Spike Score (%) 0.0222 0.0000 0.1111 0.0000 0.0111 0.0333 0.0889 0.0000
In addition to the clipping-based baselines discussed in the main text, we also compare AdaGC with several recent methods that aim to improve the stability and generalization of large language model (LLM) training, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024). The detailed results under the zero-shot setting and spike score are summarized in Table 12 and 13. The training dynamics are shown in Figures 8 and 9.
Among these methods, SPAM is designed to stabilize training by adjusting the optimizer’s behavior, while Scaled Embed and WeSaR focus on initialization or embedding scaling strategies to suppress loss spikes. Our experiments show that, although some of these methods can partly mitigate instability or improve certain metrics, AdaGC generally achieves higher stability and better final performance across model scales. Notably, while WeSaR is also effective at suppressing loss spikes, its reliance on special parameter initialization limits its applicability to from-scratch training. In contrast, AdaGC works reliably under both from-scratch and resumed training regimes, providing stronger flexibility. Overall, these results demonstrate AdaGC’s superior robustness and generalization compared to other non-clipping baselines.
Appendix E More Visualization Results
E.1 Training Dynamics
Figure 7: Llama-2 1.3B training dynamics of clipping methods.
Figure 8: Llama-2 1.3B training dynamics of other baseline methods.
Figure 9: Llama-2 7B training dynamics of other baseline methods.
E.2 Optimizer State Dynamics
Figure 10: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 11: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_3_self_attention_query_key_value of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 12: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the LMHead of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 13: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_final_layernorm of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 14: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_0_self_attention_query_key_value of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 15: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_23_input_layernorm of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Appendix F Convergence Analysis and Proof
Any operation that modifies gradients may potentially result in non-convergence. In this section, rather than providing a theoretical guarantee that AdaGC eliminates loss spikes, we present the convergence guarantee for Adam with AdaGC, stated as follows:
Theorem F.1.